The following data represent the miles per gallon for a particular make and model car for six randomly selected vehicles. Compute the​ mean, median, and mode miles per gallon.

24.1, 22.7, 37.8, 39.3, 39.1, 36.2

  

Compute the mean miles per gallon. Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.

A.The mean mileage per gallon is

nothing.

​(Round to two decimal places as​ needed.)

B.

The mean does not exist.
Answer #1

Arranging in ascending order :

22.7 , 24.1, 36.2 , 37.8 , 39.1 , 39.3

Mean = \sumXi/n = (22.7+24.1+36.2+37.8+39.1+39.3)/6 = 199.2/6 = 33.2

Mode= 39.3(highest value)

Median = ((n/2)th + (n/2 +1)th)observation/2

=(3rd +4th )obs/2= (36.2+37.8)/2 = 74/2 = 37